# Derail Yourself: Multi-turn LLM Jailbreak Attack through self-discovered clues

### v1

@@ 3.2 In-attack: first reason then attack
3. Dynamically modify the initial attack path for various victim models. During the interactions with the victim model, we propose to dynamically modify the initial attack paths to mitigate the possible misalignment between the predicted and realistic responses. We identify two typical misalignment cases and design a GPT4-Judge to assess every response from the victim model: (1) Unknown, where the victim model does not know the answer to the current query, (2) Rejective, where the victim model refuses to answer the current query. As for Unknown, we drop the attack clue, and sample another one to restart our attack again (Fig. 3 (c)), while for Rejective, we perform the toxicity reduction by removing the harmful words and using ellipsis to bypass the safety guardrails of LLMs.
0:  plain harmful query $x$, attacker model $A_{\theta}$, victim model $V_{\theta}$, iterations $N$, number of actors $K$, judge model $J_{\theta}$, monitor model $M_{\theta}$
1:  $C\leftarrow$find__\__attack__\__clues$(x,A_{\theta})$ // construct the network of attack clues
@@
Algorithm 1 ActorAttack
4 Experiments
4.1 Experimental Setup
Models. We validate the efficacy of ActorAttack on 5 prevalent LLMs: GPT-3.5 (GPT-3.5 Turbo 1106) (OpenAI, 2023), GPT-4o (OpenAI, 2024a), Claude-3.5 (Claude-3.5-sonnet-20240620) (Anthropic, 2024), Llama-3-8B (Llama-3-8B-Instruct) (Dubey et al., 2024) and Llama-3-70B (Llama-3-70B-Instruct) (Dubey et al., 2024).
Datasets. We evaluate the harmfulness of models on HarmBench (Mazeika et al., 2024), a framework that includes a harmful behaviors dataset and a wide range of both black-box and white-box attacks. For the ablation study, we uniformly sample 50 instances of HarmBench across various harmful categories as the benchmark data, and we manually check to minimize duplicates.
Attack Baselines. We compare our proposed method against several prominent attack methods, the details of baseline settings are given in App. A.1:
  1. 1.
    GCG (Zou et al., 2023b): A white-box attack that crafts adversarial examples via greedy and gradient-based discrete optimization.
  2. 2.
    PAIR (Chao et al., 2024): A black-box attack that uses an attacker LLM to automatically generate adversarial inputs for a target LLM.
  3. 3.
    PAP (Zeng et al., 2024): A black-box attack that regards LLMs as human-like communicators and persuades LLMs to jailbreak them.
  4. 4.
    CipherChat (Yuan et al., 2024b): A black-box attack that converts inputs into ciphers to jailbreak LLMs.
  5. 5.
    CodeAttack (Ren et al., 2024): A black-box attack that disguises malicious questions as a code completion task and generates harmful responses while completing the code.
  6. 6.
    Crescendo (Russinovich et al., 2024): A multi-turn black-box attack that gradually steers benign initial queries towards more harmful topics.
Attack Evaluation. To evaluate whether the attack is successful, we utilize Attack Success Rate (ASR) as our evaluation metric, which is the percentage of harmful responses given harmful queries. Following the work of (Qi et al., 2023), we utilize the robust evaluation capability of GPT-4o to provide the assessment by feeding harmful targets and model responses into the GPT-4o Judge. The judge score ranges from 1 to 5, and the higher the score is, the more harmful and more detailed the model’s responses are. We only consider an attack successful when the GPT-4o Judge assigns a score of 5. Refer to Qi et al. (2023) for details of the rubric.
Diversity Evaluation. To measure the diversity of the generated prompts across different trials, we follow the practices recommended in  (Tevet & Berant, 2020; Hong et al., 2024; Lee et al., 2024), and employ the sentence embedding distances as our metric. We embed the generated prompts using MiniLMv2 (Wang et al., 2020) encoder. We group prompts by harmful targets, and divide each group into subsets based on which turn the prompt is in. We compute the average pairwise cosine distance between the embedding of the prompts for each subset and finally average values of all subsets. See the details of the diversity metric in App. B.
Implementation Details. We configured the temperature of our attacker LLM to 1 and the temperature of the victim LLM to 0. For each harmful target, unless explicitly stated in the ablation study, ActorAttack selects 3 actors to generate 3 different multi-turn attacks, and the maximum number of queries in a multi-turn attack is set to 5.
4.2 Main Results: Effective and diverse multi-turn attack
Method Attack Success Rate($\uparrow$%)
GPT-3.5 GPT-4o Claude-3.5 Llama-3-8B Llama-3-70B Avg
single-turn GCG 55.8 12.5 3.0 34.5 17.0 24.56
PAIR 41.0 39.0 3.0 18.7 36.0 27.54
PAP 40.0 42.0 2.0 16.0 16.0 23.2
CipherChat 44.5 10.0 6.5 0 1.5 12.5
CodeAttack 67.0 70.5 39.5 46.0 66.0 57.8
multi-turn (ours) ActorAttack (w/o DM) 74.5 80.5 54.5 68.0 75.0 70.5
ActorAttack 78.5 84.5 66.5 79.0 85.5 78.8
Table 1: Attack success rate of baseline single-turn attacks and our ActorAttack on Harmbench. We present the results of ActorAttack without dynamic modification (w/o DM) for comparison with the full ActorAttack method.
ActorAttack generates more effective prompts than single-turn baselines. Table 1 shows the baseline comparison results. Although our ActorAttack method does not use any special optimization, we find that ActorAttack is the only method that achieves a high attack success rate across all target LLMs, highlighting the common and significant safety risks in the multi-turn dialogue scenario. Among the baselines, CodeAttack achieves the best performance, while its jailbreak template is hand-crafted and contains identifiable malicious instructions, making it easy to defend.
For qualitative evaluation, we provide various examples of ActorAttack, showcasing different types of human and non-human actors such as regulation, facilitation, and execution across different harmful categories, as shown in Fig. 10, Fig. 11, Fig. 12, Fig. 13 and Fig. 14. We truncate our examples to include only partial harmful information to prevent real-world harm.
ActorAttack dynamically modifies the attack path for various target models, enhancing its effectiveness. We compare the performance of our method with and without dynamic modification. As shown in Table 1, we have two findings: (1) When our attack does not involve dynamic modification and does not leverage information from the target model, our attack still exhibits good performance across different models. This indicates that our method is efficient at identifying common safety vulnerabilities of these models without requiring special adaptations. (2) The introduction of dynamic modification further improves the effectiveness of our attack by adaptively modifying the queries based on the responses from the target model, toward a more comprehensive evaluation of the safety mechanisms behind different models.
ActorAttack is also more effective and diverse than multi-turn baselines. To demonstrate the diversity and effectiveness advantages of our method, we select the most relevant multi-turn baseline, Crescendo (Russinovich et al., 2024) for comparison. To measure diversity, we run 3 independent iterations for every attack generated by each method. Fig. 4 shows the attack success rate and cosine distance between the embeddings of prompts generated by each method for different attack budgets against GPT-4o and Claude-3.5-sonnet. We find that ActorAttack is consistently more diverse, and more effective than Crescendo across various attack budgets. On the one hand, Crescendo depends on the seed instances such that its attacks could collapse to similar patterns, lacking diversity. On the other hand, Crescendo does not have an explicit reasoning chain to ensure the alignment of its queries with the harmful target, limiting its effectiveness. Qualitative assessment of examples generated by ActorAttack and Crescendo included in Fig. 7 and Fig. 8 support our analyses.
Figure 4: Attack success rate of ActorAttack and Crescendo under different attack budgets, i.e., number of queries, as well as the pairwise cosine similarity between prompts generated by ActorAttack and Crescendo for (a) GPT-4o and (b) Claude-3.5-sonnet.
4.3 Empirical Analysis
The diverse attack paths uncovered by ActorAttack are mostly effective. ActorAttack generates diverse attack paths for the same harmful target. We assess the effectiveness of every path using the score given by our judge model, and we calculate the proportion of different scores for our attack paths. As shown in Fig. 6 (a), we find that most of these paths are classified as most harmful with a top score of 5. This reveals that ActorAttack can effectively identify more safety vulnerabilities of models through its diverse attack paths.
ActorAttack finds higher-quality attacks from more diverse attack paths. One potential advantage of generating diverse attack prompts is that we can find more optimal attack paths, leading to answers of higher quality. To study this empirically, we sample different numbers of attack clues to generate diverse attacks for the same harmful target and record the best score of the attacks by our judge model. As shown in Fig. 5, we find that the proportion of attacks with a score of 5 increases with more actors (attack clues), which indicates that ActorAttack can discover more optimal attack paths by exploiting diverse attack clues.

### v2

@@ 3.2 In-attack
3. Dynamically modify the initial attack path for various victim models (optional). During the interactions with the victim model, we propose to dynamically modify the initial attack paths to mitigate the possible misalignment between the predicted and realistic responses. We identify two typical misalignment cases and design a GPT4-Judge to assess every response from the victim model: (1) Unknown failure: This occurs when the victim model responds with statements like “I don’t know how to answer this query.” To handle this, we immediately halt the current attack attempt and restart with a new one, as illustrated in Fig. 3 (c). This approach prevents unnecessary continuation of failed attack paths, thereby improving the overall efficiency of the attack process. We observed that this type of failure typically occurs during the first query, allowing us to terminate early and avoid resource wastage. (2) Rejective failure: This occurs when the victim model explicitly refuses to answer a query. In this case, we reduce the harmfulness of the query by using ellipsis to avoid sensitive words explicitly flagged by the victim model. Specifically, we refine the query by leveraging the model’s prior responses to craft follow-up queries. For instance, if the initial query mentions "bombing campaign" and the victim model flags it, we refine the query to "campaign activities" while maintaining the semantic intent. The modified query is thus more likely to bypass safety guardrails. We note that dynamic modification is proposed to enhance effectiveness, while this module is optional and our initial attacks achieve a high attack success rate as demonstrated by Table 4.
4 Experiments
4.1 Experimental Setup
Method Attack Success Rate ($\uparrow$%)
GPT-3.5 GPT-4o GPT-o1 Claude-3.5 Llama-3-8B Llama-3-70B Avg
Single-turn Attacks GCG 55.8 12.5 0.0 3.0 34.5 17.0 20.47
Multilingual 64.0 0.0 0.0 0.0 0.0 0.0 10.67
CipherChat 44.5 10.0 0.0 6.5 0 1.5 10.42
AutoDAN - - - - 37.5 38.5 38.0
PAIR 41.0 39.0 0.0 3.0 18.7 36.0 22.95
PAP 40.0 42.0 0.0 2.0 16.0 16.0 19.33
CodeAttack 67.0 70.5 2.0 39.5 46.0 66.0 48.5
ReNeLLM 76.0 69.5 12.0 55.0 68.0 24.5 50.8
Multi-turn Attacks CoA 25.5 18.8 8.0 15.5 25.5 22.5 19.3
Crescendo 60.0 62.0 14.0 38.0 60.0 62.0 49.3
ActorBreaker (ours) 78.5 84.5 60.0 78.5 79.0 85.5 77.7
Table 1: Attack success rate of single-turn attacks, multi-turn attacks and our ActorBreaker against several open and closed source LLMs on Harmbench.
Model Type Creation Execution Distribution Reception Facilitation Regulation
GPT-3.5-turbo 54% 62% 72% 54% 68% 44%
GPT-4o 44% 50% 52% 42% 44% 32%
Llama-3-8B-instruct 46% 44% 66% 40% 60% 34%
Llama-3-70B-instruct 54% 46% 62% 54% 68% 48%
Table 2: Attack success rate of different actor types of our ActorBreaker on Harmbench.
Models. We validate the efficacy of ActorBreaker on 6 prevalent LLMs: GPT-3.5 (GPT-3.5 Turbo 1106) (OpenAI, 2023), GPT-4o (OpenAI, 2024a), GPT-o1 (GPT-o1-preview) (OpenAI, 2024c), Claude-3.5 (Claude-3.5-sonnet-20240620) (Anthropic, 2024), Llama-3-8B (Llama-3-8B-Instruct) (Dubey et al., 2024) and Llama-3-70B (Llama-3-70B-Instruct) (Dubey et al., 2024).
Datasets. We evaluate all attacks on HarmBench (Mazeika et al., 2024), a framework that includes a harmful behaviors dataset and a wide range of both black-box and white-box attacks.
Attack Baselines. We compare our method against the leading attack methods on both HarmBench (Mazeika et al., 2024) and EasyJailbreak (Zhou et al., 2024b) leaderboard: GCG (Zou et al., 2023b), PAIR (Chao et al., 2024), AutoDAN (Liu et al., 2023), Multilingual (Deng et al., 2024), PAP (Zeng et al., 2024), CipherChat (Yuan et al., 2024b), CodeAttack (Ren et al., 2024), and ReNeLLM (Ding et al., 2023b), and we also select two multi-turn attack methods: CoA (Yang et al., 2024) and Crescendo (Russinovich et al., 2024). Find further details in App. D.1.
Judge Selection. We utilize Attack Success Rate (ASR) as our evaluation metric, which is the percentage of harmful responses given harmful queries. Following the work of (Qi et al., 2023; Zeng et al., 2024; Ren et al., 2024), we utilize the robust evaluation capability of GPT-4o to provide the assessment. Qi et al. (2023) shows the effectiveness and accuracy of the GPT-4 judge in identifying harmful outputs. Our human studies further confirm that the GPT-4o judge has a higher agreement with human majority voting than alternatives like Llama-Guard 2 (Team, 2024) and the OpenAI Moderation API (Markov et al., 2022). Find results in App. D.3.
Diversity Evaluation. To measure the diversity of the generated prompts across different trials, we follow the established practices in  (Tevet and Berant, 2020; Hong et al., 2024; Lee et al., 2024), and employ the BERT-sentence embedding distances as our metric. Specifically, we compute the pairwise cosine similarity between attack prompts generated across multiple trials as a measure of diversity. Further details are available in App. E.
Implementation Details. We set the temperature of our attack LLM to 1 and the victim LLM to 0. For each harmful target, unless explicitly stated in the ablation study, we select 3 actors to generate 3 multi-turn attacks, and the maximum number of queries in a multi-turn attack is set to 5. We use GPT-4o as our attack model.
4.2 Discussion of Results
ActorBreaker achieves higher ASR rates than both leading single-turn and multi-turn attack methods. Table 1 shows the baseline comparison results. Although our ActorBreaker does not use any special optimization, we find that ActorBreaker achieves the highest attack success rate across all target LLMs over both single-turn and multi-turn baselines: our attack achieves the average ASR of 77.7% as against 18.3% and 45.0% for CoA and Crescendo respectively. Such large performance gap reveals the difference of our prompt distribution with others, and demonstrates the brittleness of current LLMs to our benign yet semantically related prompts.
Further, our attack prompts are significantly more robust against GPT-o1 with strong reasoning capabilities: 60.0% for our method while 14.0% is the highest ASR among other baselines. We observe the conflicting behavior of GPT-o1 against our attacks: in its chain of thought (CoT), it first shows its safe thoughts about following the OpenAI content policies but then lists how to fulfill our query step by step (see Fig. 9). Our attack result thus raises the faithfulness concern of CoT reasoning (Lyu et al., 2023), (Lanham et al., 2023).
For qualitative evaluation, we provide various examples of ActorBreaker, showcasing the effectiveness of different types of human and non-human actors across different harmful categories (Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13). We truncate our examples to include only partial harmful information to prevent real-world harm.
The effectiveness of different types of actors The six actor categories enable the generation of diverse attack prompts, which are essential for a comprehensive probing of model safety vulnerabilities. We first validate the effectiveness of these categories. We demonstrated that prompts generated from each actor type can effectively probe vulnerabilities in LLMs. Specifically, for each harmful query, we sampled three specific actors from each category and generated multi-turn attack prompts. Table 2 showed that attack prompts from all six categories were relatively effective across both open-source and closed-source language models. This confirms the validity of our actor definitions.
C C+E C+E+D C+E+D+Rec C+E+D+Rec+F C+E+D+Rec+F+Reg
GPT-3.5-turbo 54% 66% 74% 78% 80% 84%
GPT-4o 44% 58% 68% 74% 78% 80%
Llama-3-8B-instruct 46% 60% 78% 80% 86% 86%
Llama-3-70B-instruct 54% 62% 72% 80% 88% 90%
Table 3: Attack success rate of different actor combinations of our ActorBreaker on Harmbench. The abbreviations correspond to specific actor types: Creation (C), Execution (E), Distribution (D), Reception (Rec), Facilitation (F), and Regulation (Reg).
Our six types of actors ensure diversity and comprehensive coverage of safety vulnerabilities. By using a more diverse set of actors, we can generate a wider variety of attack prompts, uncovering more safety vulnerabilities. To empirically show this, we evaluate the performance of our attacks using a different number of actor types. Results, shown in the table 3, showed that as the number of actor types increased, the overall attack success rate also improved. This highlights that different actor types target different aspects of model vulnerabilities, proving the necessity of our categorization for a comprehensive probing of model safety.
Ablation on dynamic modification. One potential advantage of our attacks is transferability. Since LLMs are pre-trained on similar web-filtered data, and possibly know the semantic associations between our prompts and the harmful target, we thus argue that the failure mode found by our attacks of one LLM might by shared by other models. To study this empirically, we compare the performance of our method with and without dynamic modification (DM). Table 4 shows that our method without DM transfers well across different LLMs and achieves the average ASR of 72.7% as against 81.2% for with DM across several aligned LLMs, demonstrating the transferability property of our method. By contrast, current multi-turn attacks like CoA and Crescendo rely on responses of the target LLM to craft their attacks, limiting their efficiency.
Higher diversity of our prompts over multi-turn baselines. To measure diversity, we run 3 independent trails for every harmful target for each method. Table 5 shows the cosine similarity between the embeddings of prompts generated by each method across various aligned LLMs. We find that ActorBreaker consistently generates more diverse prompts than CoA and Crescendo. It aligns with our analyses about attack prompts generated by CoA and Crescendo could collapse to similar patterns due to their fixed strategies and potential biases towards their seed instances, while we, grounded in social theory, can characterize the diverse semantic relationships about the toxic prompt. Qualitative assessment of examples included in Fig. 7 and Fig. 8 further supports our analyses.
Model ActorBreaker +DM
GPT-3.5 74.5 78.5
GPT-4o 80.5 84.5
Claude-3.5 65.5 78.5
Llama-3-8B 68.0 79.0
Llama-3-70B 75.0 85.5
Average 72.7 81.2
Table 4: Attack success rate of our ActorBreaker on Harmbench. We present the results of ActorBreaker with and without dynamic modification (DM).
Method GPT-4o Claude-3.5 Llama-3 Llama-3
-8B -70B
CoA 0.16 0.17 0.20 0.21
Crescendo 0.23 0.22 0.25 0.23
ActorBreaker 0.32 0.36 0.36 0.34
Table 5: The diversity of prompts generated by multi-turn attack methods. Higher values mean greater diversity. We computed the pairwise cosine similarity between attack prompts generated across multiple trials as a measure of diversity.
Figure 4: The proportion of judge scores for attacks generated by ActorBreaker, for various numbers of actors, against (a) GPT-4o and (b) Claude-3.5-sonnet. Higher score means more harmful model responses and a score of 5 means the success of the attack; (c): attack success rate of ActorBreaker against varying numbers of actors for GPT-4o and Claude-3.5-sonnet.
Higher-quality attacks result from our more diverse attack prompts. We argue that our diverse attack prompts enable us to do a wider exploration and thus find more optimal attack paths, leading to more harmful responses. To demonstrate this, given a toxic query, we sample different numbers of actors to generate multiple attacks and record the best score of the attacks by our judge model. As shown in Fig. 4, we find that the proportion of attacks with a score of 5 increases with more actors (attack clues), which indicates that ActorBreaker can discover more optimal attack paths by exploiting diverse attack clues. Find results of Llama-3-8B and Llama-3-70B in Fig. 6 of App. C.1.
Higher efficiency over multi-turn baselines. We compare our method with CoA and Crescendo in terms of time cost. Due to differences in backend technologies (e.g., vLLM (Kwon et al., 2023) or Torch (Paszke et al., 2019)) and parallelization strategies used by these methods, a fair time cost comparison is challenging. We thus propose using the average number of interactions with the target model per attack as a more consistent efficiency metric. Each turn of the attacks, including random trials, counts as one interaction. Results in Table 6 demonstrate that our method generally requires far fewer interactions to succeed compared to these baselines. Specifically, our approach achieves a 26% improvement in attack efficiency over Crescendo, confirming the efficiency advantages of our method.
Ablation on attack budgets. We evaluate the performance of our attacks under different attack budgets. We set the maximum number of conversation turns per our attack from 2 to 5. Table 7 shows that aligned LLMs are more vulnerable to our attacks in longer conversations. We argue that more number of interactions with the target LLM expands the action space of our attacks, making it more likely to find a successful attack path.
Our attack prompts bypass the toxicity detection of LLM-based input safeguard. To evaluate the harmfulness of our prompts, we employ Llama Guard 2 (Team, 2024) to classify both the original plain harmful queries and the multi-turn queries generated by our attack and other multi-turn attacks to be safe or unsafe. The classifier score represents the probability of being “unsafe.” Fig. 5 (a) shows that the toxicity of our multi-turn queries is much lower than that of both the original harmful query and the queries generated by Crescendo in both GPT-4o and Claude-3.5-sonnet. We note that the toxicity of prompts generated by CoA becomes lower than both our attack and Crescendo with more attack turns. This is because we observe that the prompts of CoA gradually deviate from the harmful target with its further interactions with the target LLM. Therefore, though being less harmful, CoA is less effective than our attack and Crescendo, as shown in Table 1.
Model CoA Crescendo ActorBreaker
GPT-3.5 15.8 12.0 8.5
GPT-4o 14.6 11.5 8.1
Claude-3.5 43.3 14.9 10.9
Llama-3-8B 14.6 10.5 8.3
Llama-3-70B 13.6 10.3 8.0
Avg 20.4 11.8 8.7
Table 6: Time cost of multi-turn attacks on Harmbench. We select the average number of queries per harmful target as the proxy of time cost.
Model Number of queries
2 3 4 5
GPT-4o 51.0 65.0 70.0 84.5
Claude-3.5 41.5 53.0 65.0 78.5
Llama-3-8B 67.0 72.0 77.0 79.0
Llama-3-70B 74.0 79.0 84.5 85.5
Table 7: Attack success rate of ActorBreaker on Harmbench within different attack budgets. We set the maximum number of conversation turns per our multi-turn attack from 2 to 5.
Adaptive defense: multi-turn safety data construction. Since current safety alignment datasets (Ji et al., 2024; Bai et al., 2022) mainly focus on single-turn Q-A pairs, we thus propose to construct a multi-turn safety dataset using our attack prompts to mitigate the safety gap. We propose to use the judge model to detect where the victim model first elicits harmful responses in the multi-turn conversations and insert the refusal responses here. For the training data, we sample 600 harmful instructions from Circuit Breaker (Zou et al., 2024b), which have been filtered to avoid data contamination with the Harmbench and construct 1680 multi-turn safety prompts. Further details can be found in App. D.2.
Stability of our attacks against existing defenses and our adaptive defense. Besides our adaptive defense, we also select three distinct and state-of-the-art defense baselines: Rephrase (Jain et al., 2023), RPO (Zhou et al., 2024a), and Circuit Breaker (CB) (Zou et al., 2024a), to comprehensively assess the effectiveness of our attack. Since both Circuit Breaker and our defense mechanism rely on fine-tuning, we report the results specifically for Llama-3-8B-Instruct. Table 8 show that Rephrase and RPO offer a partial reduction in ASR. It demonstrates that our attacks are robust against semantically meaningful or random perturbations due to naturalness of our prompts without extra optimization or specialized techniques. However, Circuit Breaker greatly reduces the success rate of our attack, demonstrating the potential of safety alignment within the representation space. Moreover, we find that the CB model trained on our multi-turn dataset demonstrates greater robustness against multi-turn attacks compared to CB trained on single-turn data. This highlights the value of our multi-turn safety data. We also found that CB trained on our multi-turn data is more robust than SFT trained on the same dataset, highlighting the algorithmic advantages of CB over SFT. The helpfulness evaluation results of our fine-tuned model are in App. D.2.
5 Conclusion
This paper highlights a critical blind spot in the safety mechanisms of aligned LLMs: their vulnerability to natural distribution shifts. We discovered that seemingly innocent prompts, which are semantically linked to harmful content, can bypass current safety mechanisms and lead to unsafe model behavior. Our proposed solution, ActorBreaker, offers a novel way to systematically probe LLMs for these vulnerabilities using multi-turn prompts grounded in actor-network theory. Experimental results confirm that ActorBreaker achieves superior performance compared to other attack methods. To mitigate these risks, we emphasize the importance of expanding safety training data to address the broader semantic landscape of toxic content. Fine-tuning LLMs on our dataset generated by ActorBreaker greatly improves robustness against such attacks.
(a) Toxicity of prompts against GPT-4o
(b) Toxicity of prompts against Claude-3.5
Figure 5: The classifier score produced by LlamaGuard 2 for both plain harmful queries and multi-turn attack queries against GPT-4o (a) and Claude-3.5-sonnet (b). The classifier score represents the probability of being “unsafe” of the prompt.
Method Llama-3-8B GPT-3.5 GPT-4o
No Defense 78.0 78.5 84.5
Rephrase 54.0 50.0 80.0
RPO 54.0 42.0 50.0
CB + single-turn data 28.0 - -
CB + our multi-turn data 16.5 - -
SFT + our multi-turn data 32.0 - -
Table 8: Attack success rate (%) of our ActorBreaker against various defense methods. "SFT" indicates supervised fine-tuning using safety data and "CB" denotes Circuit Breaker; multi-turn data includes context-aware adversarial prompts.
6 Limitations
In this study, we focus on generating actors related to harmful targets in English, without considering multilingual scenarios. Different languages come with distinct cultures and histories, which means that for the same harmful behavior, actors associated with different languages may differ. Since LLMs have demonstrated strong multilingual capabilities (Nguyen et al., 2023; Sengupta et al., 2023; Workshop et al., 2022), it would be valuable to study our attack methods across multiple languages for better coverage of the real-world distribution of actors. Future work can also explore the applicability of our method to jailbreak multi-modal models (Liu et al., 2024c, b). For defense, we use safety fine-tuning to generate refusal responses. However, we observe a trade-off between helpfulness and safety. Exploring reinforcement learning from human feedback (RLHF) in the multi-turn dialogue scenarios could be a valuable direction, e.g., designing a reward model that provides more granular scoring at each step of multi-turn dialogues.
7 Ethics Statement
We propose an automated method to generate jailbreak prompts for multi-turn dialogues, which could potentially be misused to attack commercial LLMs. However, since multi-turn dialogues are a typical interaction scenario between users and LLMs, we believe it is necessary to study the risks involved to better mitigate these vulnerabilities. We followed ethical guidelines throughout our study. To minimize real-world harm, we will disclose the results to major LLM developers before publication. Additionally, we explored using data generated by ActorBreaker for safety fine-tuning to mitigate the risks. We commit to continuously monitoring and updating our research in line with technological advancements.
Acknowledgements
This project was supported by Shanghai Artificial Intelligence Laboratory, the National Natural Science Foundation of China (Grant No. 72192821), YuCaiKe (Grant No. 231111310300), the Fundamental Research Funds for the Central Universities (Grant No. YG2023QNA35), and the National Natural Science Foundation of China (Grant No. 62472282).
Appendix A Related Work
Defenses for LLMs. To ensure LLMs safely follow human intents, various defense measures have been developed, including prompt engineering (Xie et al., 2023; Zheng et al., 2024), aligning models with human values (Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024; Meng et al., 2024; Yuan et al., 2024a), model unlearning (Li et al., 2024b; Zhang et al., 2024b), representation engineering(Zou et al., 2024b) and implementing input and output guardrails (Dubey et al., 2024; Inan et al., 2023; Zou et al., 2024a). Specifically, input and output guardrails involve input perturbation (Robey et al., 2023; Cao et al., 2023; Liu et al., 2024e), safety decoding (Xu et al., 2024), and jailbreak detection (Zhang et al., 2024c; Yuan et al., 2024c; Phute et al., 2023; Alon and Kamfonas, 2023; Jain et al., 2023; Hu et al., 2024). Priority training also shows its effectiveness by training LLMs to prioritize safe instructions (Lu et al., 2024; Wallace et al., 2024; Zhang et al., 2023).
Appendix B Algorithm
Notations for Algorithm 1. Except for the victim model, we use the same LLM to implement the other three models via different instructions. $H$ denotes the history of the dialogue and $C_{retry}$ represents the number of attempts currently made.
0:  A toxic query $x$, Attack model $A_{\theta}$ that generates multi-turn prompts, Victim model being attacked $V_{\theta}$, Judge model $J_{\theta}$ that determines the success of attacks, Monitor model $M_{\theta}$ (optional) that decides whether to modify the current prompt, Iterations $N$, Number of actors $K$
// construct the network of attack clues
1:  $C\leftarrow$find__\__attack__\__clues$(x,A_{\theta})$
@@ Appendix B Algorithm
Algorithm 1 ActorBreaker
Appendix C Additional results
C.1 Additional results for ablation on the number of actors.